OpenStudy (anonymous):

write the equation in slope intercept form of the line that has a slope o2 and contains points (4,2)

pooja195 (pooja195):

o2?

OpenStudy (anonymous):

I meant to say of 2 sorry

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\] m=slope (x1,y1)

pooja195 (pooja195):

plug in the numbers to the point slope formula i put above

OpenStudy (anonymous):

so is it y-2=2(x-4)

pooja195 (pooja195):

yes now solve

pooja195 (pooja195):

1) Distribute \[\huge~2(x-4) \]

OpenStudy (anonymous):

2x-8

pooja195 (pooja195):

good so y-2=2x-8 -8+2=?

OpenStudy (anonymous):

-6

pooja195 (pooja195):

y=2x-6 would be the final answer then :)